OpenStudy (anonymous):

Complete combustion of 8.80 g of a hydrocarbon produced 26.7 g of CO2 and 13.6 g of H2O. What is the empirical formula for the hydrocarbon?

Parth (parthkohli):

\[\rm hydrocarbon + CO_2 \longrightarrow H_2O + CO_2\]Number of moles of \(\rm H_2O = 13.6/18\). Number of moles of \(\rm H = 13.6/ 9\). Number of moles of \(\rm CO_2 = 26.7/44\). Number of moles of \(\rm C = 26.7/44\). The ratio of \(\rm H:C\) is \(13.6/9 \div 26.7/44 \approx 5:2\). The empirical formula, thus, seems to be \(\rm C_2 H_5\).

OpenStudy (koikkara):

\(\color{red}{~\Large\star~``Welcome~to~Open~Study~!!"~\star~}\) \(Hope~it~helps~!!\) Don't Forget, \(Appreciate~the~Best~Answer\) with a click on \(\color{purple}{"Best Response"}\) Button Too ! \(\color{darkgreen}{~''Nice~To~Meet~You~!!"}\) @ rperedo

OpenStudy (koikkara):

@rperedo